OpenStudy (anonymous):

Each side of a equilateral triangle measures 9cm. Find the height, h, of the triangle. ...Please show work :] thanks

OpenStudy (anonymous):

break it up into 2 triangles. base is 4.5, hypotenuse is 9 4.5^2+b^2=81 20.25+b^2=81 b^2=60.75 h=7.794228634059948